# KKL theorem for the influence of a set of variables

### v1

@@ 3.2. Influence
The derivative neatly lays the ground for defining the classical notion of the (single-bit) influence of bit $i$ on $f$ by
$\operatorname{Inf}_{i}(f):=\sum_{\begin{subarray}{c}\ \mathbf{j}\subseteq[n]\!\colon\\ \mathbf{j}\ni i\end{subarray}}\hat{f}(\mathbf{j})^{2}=\|\partial_{i}f\|_{2}^{2}.$
We follow up with the respective multi-bit definition announced in the introduction.
Definition 3.5.
Let $f\colon\{-1,1\}^{n}\to\mathbb{R}$ and $\mathbf{i}\subseteq[n]$. We define the influence of $\mathbf{i}$ on $f$ by
$\operatorname{Inf}_{\mathbf{i}}(f):=\sum_{\begin{subarray}{c}\ \mathbf{j}\subseteq[n]\!\colon\\ \mathbf{j}\supseteq\mathbf{i}\end{subarray}}\hat{f}(\mathbf{j})^{2}=\|\partial_{\mathbf{i}}f\|_{2}^{2}.$
Approximately speaking, $\operatorname{Inf}_{\mathbf{i}}(f)$ measures if all the bits of $\mathbf{i}$ are influential on $f$. Sometimes, we will say that $\operatorname{Inf}_{\mathbf{i}}(f)$ is a $d$-degree influence, where $d=|\mathbf{i}|$. The above definition was introduced by Tal [11] in his study of the Fourier mass at high levels in $\mathbb{AC^{0}}$ Boolean circuits.
Several other generalizations of the single-bit influence have been studied, see [2]. Probably the most popular is that of Ben-Or and Linial [1], where for $f\colon\{-1,1\}^{n}\to\{-1,1\}$ and $\mathbf{i}\subseteq[n]$ they define
$\mathcal{I}_{\mathbf{i}}(f):\!\!$ $=\underset{\mathbf{x}\sim\mu_{n}}{\operatorname{\mathbb{P}}}(\exists y_{\mathbf{x}}\in\{-1,1\}^{n}\colon{y_{\mathbf{x}}|}_{[n]\setminus\mathbf{i}}={\mathbf{x}|}_{[n]\setminus\mathbf{i}}\text{ and }f(\mathbf{x})\neq f(y_{\mathbf{x}}))$ (4)
$=\underset{\mathbf{x}\sim\mu_{n}}{\operatorname{\mathbb{P}}}(f\text{ is not fully determined by bits on positions }\mathbf{i}\text{ of }\mathbf{x}).$
In a way, Definition 3.5 complements the Ben-Or and Linial definition: Definition 3.5 (as we will see) has better analytic properties than (4). At the same time it lacks a precise non-analytic description, which (4) has.
Intuitively, Definition (4) measures if any of the bits in $\mathbf{i}$ is influential, while Definition 3.5 measures if all the bits in $\mathbf{i}$ are. While the latter is not entirely true, it serves as a good rule of thumb. Indeed, Proposition 3.6 (ii) below says that $\operatorname{Inf}_{\mathbf{i}}(f)$ is proportional to the measure of the support of $\partial_{\mathbf{i}}f$, which by Proposition 3.4 (roughly) corresponds to every bit of $\mathbf{i}$ being influential. Further, we have $\mathcal{I}_{\mathbf{i}}(f)\geqslant\operatorname{Inf}_{\mathbf{i}}(f)$.
@@
Observe that by Proposition 3.4 we have
$\{x\in\{-1,1\}^{n}\colon f\text{ is not fully determined by bits }\mathbf{i}\text{ of }x\}\supseteq\{x\in\{-1,1\}^{n}\colon\partial_{\mathbf{i}}f\neq 0\},$
so $\mathcal{I}_{\mathbf{i}}(f)\geqslant\operatorname{\mathbb{P}}(\partial_{\mathbf{i}}f\neq 0)$.
Further, by Corollary 3.3 we have $(\partial_{\mathbf{i}}f)^{2}\in[\operatorname{\mathbbm{1}}_{\{\partial_{\mathbf{i}}f\neq 0\}}/2^{2r-2},\operatorname{\mathbbm{1}}_{\{\partial_{\mathbf{i}}f\neq 0\}}]$, so taking expectation yields (ii) and (i).
For (iii), again, by Corollary 3.3 we obtain
$\operatorname{Inf}_{\mathbf{i}}(f)=\operatorname{\mathbb{E}}[|\partial_{\mathbf{i}}f|^{2}]\geqslant\operatorname{\mathbb{E}}\Big{[}\frac{1}{2^{r-1}}|\partial_{\mathbf{i}}f|\Big{]}\geqslant\frac{1}{2^{r-1}}|\operatorname{\mathbb{E}}[\partial_{\mathbf{i}}f]|=\frac{1}{2^{r-1}}|\hat{f}(\mathbf{i})|.\qed$
@@ 3.3. Hypercontraction and the log-Sobolev inequality
When $\operatorname{Inf}_{\mathbf{i}}(f)\in\{0,1\}$ we let $1\cdot(\ln 1/1)^{-l}=+\infty$ and $00\cdot(\ln 1/0)^{-l}=$ 0.
@@
If $\operatorname{Inf}_{\mathbf{i}}(f)=1$, then the right-hand side equals +$\infty$+ ∞. If $\operatorname{Inf}_{\mathbf{i}}(f)=0$, then $\partial_{\mathbf{i}}f\equiv 0$, so the integral is equal to 00. Therefore, in both cases the statement is true. Henceforth, we will assume that $\operatorname{Inf}_{\mathbf{i}}(f)\in(0,1)$.
@@
where we used the bounds $2^{(2d-l-3)u}\leqslant 2^{2d-l-1}$ and $2^{u}\leqslant 1+u$ for $u\in[0,1]$.
@@ 5. The hypertribe example
In the case of the KKL theorem, the sharpness is provided by the (standard) tribe functions due to Ben-Or and Linial [1]. For the construction, one splits $n$ voters (bits) into disjoint tribes of sizes roughly $\log(\frac{n}{\log n})$ and declares the outcome (i.e. output of the function) to be $+1$ if there is a tribe such that all its members voted $+1$, and to be $-1$, if there is no such tribe.
@@
We start by stating Theorem 5.1, a special case of Theorem 1 from [8]. Based on this, in Example 5.2 we construct the hypertribe functions $H_{n}^{d}$. In Proposition 5.3, we state some basic quantitative observations regarding the structure of $H_{n}^{d}$. In Propositions 5.5, 5.9 and 5.8 we prove bounds on the influences of $H^{d}_{n}$ and the probabilities of the outcomes ±$\pm 1$. Finally, in Theorem 5.10 we collect the proven bounds and show that the hypertribes indeed asymptotically attain the lower bound from Theorem 1.6.
@@
[…]
Let us treat $+1$ and $-1$ as the Boolean true and false, respectively. We construct the $d$-hypertribe function $H_{n}^{d}\colon\{-1,1\}^{n}\to\{-1,1\}$ as follows
@@
    |$\big{|}\{A\in\mathcal{F}_{n}\colon\mathbf{i}\subseteq A\}\big{|}\leqslant\left(\frac{n}{\log_{2}n}\right)^{d-|\mathbf{i}|};$
@@
[…]
    |$\big{|}\{A\in\mathcal{F}_{n}\colon\mathbf{i}\subseteq A\}\big{|}\leqslant\frac{{n-s\choose d-s}}{{k-s\choose d-s}}\leqslant\frac{n^{d-s}}{k^{d-s}}\leqslant\left(\frac{n}{\log_{2}n}\right)^{d-s}.$
@@
    The upper bound follows from (P1) applied to $\mathbf{i}=\varnothing$. For the lower bound note that every tribe covers ($k\choose d$ $d$-sets. Hence, by ()($\ast$)( ∗ ),
@@
There is a constant $c_{2}$ depending only on $d$ such that for every $r\in[d]$,
$\operatorname{MaxInf}_{r}(H_{n})\leqslant c_{2}\left(\frac{\log_{2}n}{n}\right)^{r}.$
@@
Fix $\mathbf{i}\in[n]_{r}$ for some $r\in[d]$. By Proposition 3.6 we have
$\operatorname{Inf}_{\mathbf{i}}(H_{n})\leqslant\operatorname{\mathbb{P}}(\partial_{\mathbf{i}}H_{n}\neq 0).$ (9)
Choose $x\in\{-1,1\}^{n}$ for which $\partial_{\mathbf{i}}H_{n}(x)\neq 0$ and fix $i\in\mathbf{i}$. In light of Proposition 3.4, there exists $\varepsilon\in\{-1,1\}^{\mathbf{i}\setminus i}$ satisfying
$H_{n}(x^{\mathbf{i}\to\varepsilon^{+}})\neq H_{n}(x^{\mathbf{i}\to\varepsilon^{-}}).$ (10)
As $H_{n}$ is increasing, (10) may be rewritten as
$H_{n}(x^{\mathbf{i}\to\varepsilon^{+}})=+1\quad\text{and}\quad H_{n}(x^{\mathbf{i}\to\varepsilon^{-}})=-1.$
It follows that there is a tribe $A\in\mathcal{F}_{n}$ (which may not be unique) such that $i\in A$ and $x^{(\mathbf{i}\setminus i)\to\varepsilon}\in\mathcal{T}(A\setminus i)$. In particular, $A\in\mathcal{F}_{n}$ is such that $i\in A$ and $x\in\mathcal{T}(A\setminus\mathbf{i})$.
@@
$\operatorname{\mathbb{P}}(\partial_{\mathbf{i}}H_{n}\neq 0)$ $\leqslant\operatorname{\mathbb{P}}\Bigg{(}\bigcup_{S\in B_{\mathbf{i}}}\bigcup_{\begin{subarray}{c}\{A_{\mathbf{j}}\}_{\mathbf{j}\in S}\subseteq\mathcal{F}_{n}\colon\\ \forall\mathbf{j}\in S\colon\mathbf{j}\subseteq A_{\mathbf{j}}\end{subarray}}\bigcap_{\mathbf{j}\in S}\mathcal{T}(A_{\mathbf{j}}\setminus\mathbf{i})\Bigg{)}$
@@
As the number of partitions of $\mathbf{i}$ is bounded by some function of $d$, by (9) and (11) it is enough to show that for every partition $S\in B_{\mathbf{i}}$ we have
@@
Clearly, $\bigcap_{\mathbf{j}\in S}\mathcal{T}(A_{\mathbf{j}}\setminus\mathbf{i})=\mathcal{T}\big{(}\bigcup_{\mathbf{j}\in S}A_{\mathbf{j}}\setminus\mathbf{i}\big{)}$. Further, as $|S|\leqslant|\mathbf{i}|\leqslant d$ and (P3) holds,
$\bigg{|}\bigcup_{\mathbf{j}\in S}A_{\mathbf{j}}\setminus\mathbf{i}\bigg{|}\geqslant|S|(k-(d-1)^{2})-d\geqslant|S|k-d^{3},$
@@
In the next two propositions, we prove that $H_{n}$ attains both values ±$\pm 1$ with at least positive constant probability. For the proof of Proposition 5.8 we use the Harris inequality for decreasing sets, which we recall below. In Proposition 5.9 the argument proceeds via a standard second moment method. Afterwards we finally prove Theorem 5.10.
@@
$\operatorname{MaxInf}_{d}(H_{n})\leqslant c_{5}\mathbf{W}^{\geqslant d}(H_{n})\left(\frac{\log_{2}n}{n}\right)^{d}.$
In particular, for every $d\geqslant 2$ Theorem 1.6 is asymptotically sharp.
@@
By Proposition 5.5 we have $\operatorname{MaxInf}_{d}(H_{n})\leqslant c_{2}\big{(}\frac{\log_{2}n}{n}\big{)}^{d}$. Hence, it is enough to show that there is a constant $c_{6}>0$ such that for sufficiently large $n$ we have $\mathbf{W}^{\geqslant d}(H_{n})>c_{6}$.
@@
$\mathbf{W}^{\geqslant d}(H_{n})=1-\widehat{H_{n}}(\varnothing)^{2}-\sum_{r=1}^{d-1}\mathbf{W}^{=r}(H_{n}).$ (16)
@@
$1-\widehat{H_{n}}(\varnothing)^{2}=4\operatorname{\mathbb{P}}(H_{n}=+1)\operatorname{\mathbb{P}}(H_{n}=-1)\geqslant 4c_{3}c_{4}.$ (17)
@@
$\mathbf{W}^{=r}(H_{n})=\sum_{\mathbf{i}\in[n]_{r}}\widehat{H_{n}}(\mathbf{i})^{2}\leqslant 2^{2d}\sum_{\mathbf{i}\in[n]_{r}}(\operatorname{Inf}_{\mathbf{i}}(H_{n}))^{2}\leqslant 2^{2d}n^{r}(\operatorname{MaxInf}_{r}(H_{n}))^{2}$
and by Proposition 5.5 this may be further bounded above by

### v2

@@ 3.2. Influence
In the following section we will connect the influence definitions from the introduction with the notion of derivative, and prove some useful bounds between them.
Firstly, let us connect the definition of influence of a single-bit with the Fourier-Walsh coefficients. Observe that $f(x)\neq f(x^{\oplus i})$ occurs if and only if $\partial_{i}f(x)\neq 0$, which occurs if and only if $|\partial_{i}f(x)|=1$. Therefore, by Parseval’s identity
$\operatorname{Inf}_{i}(f)=\operatorname{\mathbb{P}}(\partial_{i}f(\mathbf{x})\neq 0)=\operatorname{\mathbb{E}}[(\partial_{i}f(\mathbf{x}))^{2}]=\|\partial_{i}f\|_{2}^{2}=\sum_{\begin{subarray}{c}\ \mathbf{j}\subseteq[n]\!\colon\\ \mathbf{j}\ni i\end{subarray}}\hat{f}(\mathbf{j})^{2}.$
In particular, as we promised in the introduction, $\operatorname{Inf}_{i}(f)$ is equal to the T-influence $\operatorname{Inf}_{\{i\}}(f)$.
The definition of the multi-bit derivative yields an analogous representation of the T-influence. Let $f\colon\{-1,1\}^{n}\to\mathbb{R}$ and $\mathbf{i}\subseteq[n]$. Again, by Parseval’s identity we have
$\operatorname{Inf}_{\mathbf{i}}(f)=\sum_{\begin{subarray}{c}\ \mathbf{j}\subseteq[n]\!\colon\\ \mathbf{j}\supseteq\mathbf{i}\end{subarray}}\hat{f}(\mathbf{j})^{2}=\|\partial_{\mathbf{i}}f\|_{2}^{2}.$
@@
By Corollary 3.3 we have $(\partial_{\mathbf{i}}f)^{2}\in[\operatorname{\mathbbm{1}}_{\{\partial_{\mathbf{i}}f\neq 0\}}/2^{2r-2},\operatorname{\mathbbm{1}}_{\{\partial_{\mathbf{i}}f\neq 0\}}]$, so taking expectation yields (i).
For (ii), again, by Corollary 3.3 we obtain
$\operatorname{Inf}_{\mathbf{i}}(f)=\operatorname{\mathbb{E}}[|\partial_{\mathbf{i}}f|^{2}]\geqslant\operatorname{\mathbb{E}}\Big{[}\frac{1}{2^{r-1}}|\partial_{\mathbf{i}}f|\Big{]}\geqslant\frac{1}{2^{r-1}}|\operatorname{\mathbb{E}}[\partial_{\mathbf{i}}f]|=\frac{1}{2^{r-1}}|\hat{f}(\mathbf{i})|.$
Proposition 3.5.
Let $f\colon\{-1,1\}^{n}\to\{-1,1\}$ and a non-empty $\mathbf{i}\subseteq[n]$ satisfy $\partial_{\mathbf{i}}f(x)\neq 0$ for some $x\in\{-1,1\}^{n}$. Then every $i\in\mathbf{i}$ is $\mathbf{i}$-pivotal for $f$ on $x$. In particular, $\operatorname{JInf}_{\mathbf{i}}(f)\geqslant\operatorname{Inf}_{\mathbf{i}}(f)$.
Proof.
Take an arbitrary $i\in\mathbf{i}$, and denote by $y^{+}=y^{i\to+1}$ and $y^{-}=y^{i\to-1}$. By Proposition 3.2 we have
$=\frac{1}{2^{|\mathbf{i}|}}\sum_{y\in\{-1,1\}^{\mathbf{i}\setminus i}}\Big{(}f(x^{\mathbf{i}\to y^{+}})-f(x^{\mathbf{i}\to y^{-}})\Big{)}\prod_{j\in\mathbf{i}\setminus i}y_{j}.$
Since $\partial_{\mathbf{i}}f(x)\neq 0$, there exists a non-zero summand in the above sum. As a result, $i$ is $\mathbf{i}$-pivotal for $f$ on $x$, which holds for every $i\in\mathbf{i}$.
For the inequality part, the above implies $\operatorname{JInf}_{\mathbf{i}}(f)\geqslant\operatorname{\mathbb{P}}(\partial_{\mathbf{i}}f\neq 0)$ and so by Proposition 3.4 we get $\operatorname{JInf}_{\mathbf{i}}(f)\geqslant\operatorname{Inf}_{\mathbf{i}}(f)$. ∎
Remark 3.6.
For $\mathbf{i}$ of size $1$ or $2$ one can check that the first part of the proposition is actually an equivalence, in which case $\operatorname{JInf}_{\mathbf{i}}(f)=\operatorname{\mathbb{P}}(\partial_{\mathbf{i}}f\neq 0)$.
@@ 3.3. Hypercontraction and the log-Sobolev inequality
When $\operatorname{Inf}_{\mathbf{i}}(f)\in\{0,1\}$ we let $1\cdot(\ln 1/1)^{-l}=+\infty$ and $0\cdot(\ln 1/0)^{-l}=0$.
@@
If $\operatorname{Inf}_{\mathbf{i}}(f)=1$, then the right-hand side equals $\infty$. If $\operatorname{Inf}_{\mathbf{i}}(f)=0$, then $\partial_{\mathbf{i}}f\equiv 0$, so the integral is equal to 0. Therefore, in both cases the statement is true. Henceforth, we will assume that $\operatorname{Inf}_{\mathbf{i}}(f)\in(0,1)$.
@@
where we used the bounds $2^{l-1+2(d-1)u}\leqslant 2^{2d-2+(l+1)u}$ and $2^{u}\leqslant 1+u$ for $u\in[0,1]$.
@@ 5. The hypertribe example
In the case of the KKL theorem, the sharpness is provided by the (standard) tribe functions due to Ben-Or and Linial [2]. For the construction, one splits $n$ voters (bits) into disjoint tribes of sizes roughly $\log(\frac{n}{\log n})$ and declares the outcome (i.e. output of the function) to be +1+1 if there is a tribe such that all its members voted +1+1, and to be 1-1, if there is no such tribe.
@@
We start by stating Theorem 5.1, a special case of Theorem 1 from [11]. Based on this, in Example 5.2 we construct the hypertribe functions $H_{n}^{d}$. In Proposition 5.3, we state some basic quantitative observations regarding the structure of $H_{n}^{d}$. In Propositions 5.5, 5.9 and 5.8 we prove bounds on the influences of $H_{n}^{d}$ and the probabilities of the outcomes $\pm 1$. Finally, in Theorem 5.10 we collect the proven bounds and show that the hypertribes indeed asymptotically attain the lower bound from Theorem 1.5.
@@
($d$-hypertribes) Fix $d\geqslant 2$. Let $k(n)=d\log_{2}\frac{n}{\log_{2}n}$, which is an integer when $n=2^{2^{a}}$ and $a\in\mathbb{Z}_{+}$. For sufficiently large $n$ we clearly have $k(n)\leqslant\frac{1}{2}\sqrt{n}$ and $k(n)\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}\leavevmode\nobreak\ \infty$. Choose $n>N$ so that $k(n)$ is integer. Then, by Theorem 5.1, we get a family $\mathcal{F}_{n}$ of $k(n)$-element subsets of $[n]$ satisfying both ($\ast$) and ($\ast\ast$).
Let us treat +1+1 and 1-1 as the Boolean true and false, respectively. We construct the $d$-hypertribe function $H_{n}^{d}\colon\{-1,1\}^{n}\to\{-1,1\}$ as follows
@@
    $\big{|}\{A\in\mathcal{F}_{n}\colon\mathbf{i}\subseteq A\}\big{|}\leqslant\left(\frac{n}{\log_{2}n}\right)^{d-|\mathbf{i}|};$
@@
    Fix $\mathbf{i}\in[n]_{\leqslant d}$ and let $s=|\mathbf{i}|$. There are $n-s\choose d-s$ $d$-sets in $n]_{d}$ containing $\mathbf{i}$. Every tribe $A\in\mathcal{F}_{n}$ such that $\mathbf{i}\subseteq A$ covers exactly $k-s\choose d-s$ $d$-sets containing $\mathbf{i}$. Therefore, by $(\ast\ast)$,
    $\big{|}\{A\in\mathcal{F}_{n}\colon\mathbf{i}\subseteq A\}\big{|}\leqslant\frac{{n-s\choose d-s}}{{k-s\choose d-s}}\leqslant\frac{n^{d-s}}{k^{d-s}}\leqslant\left(\frac{n}{\log_{2}n}\right)^{d-s}.$
@@
    The upper bound follows from (P1) applied to $\mathbf{i}=\varnothing$. For the lower bound note that every tribe covers $k\choose d$ $d$-sets. Hence, by $(\ast)$,
@@
There is a constant $c_{2}$ depending only on $d$ such that for every non-empty $\mathbf{i}\subseteq[n]$ of size at most $d$,
$\operatorname{JInf}_{\mathbf{i}}(H_{n})\leqslant c_{2}\left(\frac{\log_{2}n}{n}\right)^{|\mathbf{i}|}.$
@@
Consider $x\in\{-1,1\}^{n}$ such that every $i\in\mathbf{i}$ is $\mathbf{i}$-pivotal for $H_{n}$ on $x$. By the definition of the $\mathbf{i}$-pivotality, for every $i$ there is $y\in\{-1,1\}^{n}$ such that $y_{j}=x_{j}$ for every $j\in[n]\setminus\mathbf{i}$ and
$H_{n}(y^{i\to+1})\neq H_{n}(y^{i\to-1}).$
As $H_{n}$ is increasing, the above may be rewritten as
$H_{n}(y^{i\to+1})=1\quad\text{and}\quad H_{n}(y^{i\to-1})=-1.$
It follows that there is a tribe $A\in\mathcal{F}_{n}$ (which may not be unique) such that $i\in A$ and $x\in\mathcal{T}(A\setminus\mathbf{i})$.
@@
$\operatorname{JInf}_{\mathbf{i}}(H_{n})$ $\leqslant\operatorname{\mathbb{P}}\Bigg{(}\bigcup_{S\in B_{\mathbf{i}}}\bigcup_{\begin{subarray}{c}\{A_{\mathbf{j}}\}_{\mathbf{j}\in S}\subseteq\mathcal{F}_{n}\colon\\ \forall\mathbf{j}\in S\colon\mathbf{j}\subseteq A_{\mathbf{j}}\end{subarray}}\bigcap_{\mathbf{j}\in S}\mathcal{T}(A_{\mathbf{j}}\setminus\mathbf{i})\Bigg{)}$
@@
As the number of partitions of $\mathbf{i}$ is bounded by some function of $d$, by (8) it is enough to show that for every partition $S\in B_{\mathbf{i}}$ we have
@@
Clearly, $\bigcap_{\mathbf{j}\in S}\mathcal{T}(A_{\mathbf{j}}\setminus\mathbf{i})=\mathcal{T}\big{(}\bigcup_{\mathbf{j}\in S}A_{\mathbf{j}}\setminus\mathbf{i}\big{)}$. Further, as $|S|\leqslant|\mathbf{i}|\leqslant d$ and (P3), by the inclusion-exclusion principle,
$\bigg{|}\bigcup_{\mathbf{j}\in S}A_{\mathbf{j}}\setminus\mathbf{i}\bigg{|}\geqslant|S|k-{|S|\choose 2}(d-1)-d\geqslant|S|k-d^{3},$
@@
In the next two propositions, we prove that $H_{n}$ attains both values $\pm 1$ with at least positive constant probability. For the proof of Proposition 5.8 we use the Harris inequality for decreasing sets, which we recall below. In Proposition 5.9 the argument proceeds via a standard second moment method. Afterwards we finally prove Theorem 5.10.
@@
$\operatorname{JInf}_{\mathbf{i}}(H_{n})\leqslant c_{5}\mathbf{W}^{\geqslant d}(H_{n})\left(\frac{\log_{2}n}{n}\right)^{d}$
for all $\mathbf{i}\in[n]^{d}$. In particular, for every $d\geqslant 2$ Theorem 1.5 is asymptotically sharp.
@@
By Proposition 5.5 we have $\operatorname{JInf}_{\mathbf{i}}(H_{n})\leqslant c_{2}\big{(}\frac{\log_{2}n}{n}\big{)}^{d}$ for all $\mathbf{i}\in[n]^{d}$. Hence, it is enough to show that there is a constant $c_{6}>0$ such that for sufficiently large $n$ we have $\mathbf{W}^{\geqslant d}(H_{n})>c_{6}$.
@@
$\mathbf{W}^{\geqslant d}(H_{n})=1-\widehat{H^{n}}(\varnothing)^{2}-\sum_{r=1}^{d-1}\mathbf{W}^{=r}(H_{n}).$ (13)
@@
$1-\widehat{H^{n}}(\varnothing)^{2}=4\operatorname{\mathbb{P}}(H_{n}=+1)\operatorname{\mathbb{P}}(H_{n}=-1)\geqslant 4c_{3}c_{4}.$ (14)
@@
$\mathbf{W}^{=r}(H_{n})=\sum_{\mathbf{i}\in[n]_{r}}\widehat{H^{n}}(\mathbf{i})^{2}\leqslant 2^{2d}\sum_{\mathbf{i}\in[n]_{r}}(\operatorname{Inf}_{\mathbf{i}}(H_{n}))^{2}\leqslant 2^{2d}n^{r}(\operatorname{MaxInf}_{r}(H_{n}))^{2}$
and by Propositions 3.5 and 5.5 this may be further bounded above by